# Consequences of a Two-Time Relativistic Bohmian Model

### v14

@@ I Introduction
    Rejecting the decomposition of the wavefunction into norm and phase, and interpreting it as an real quantum field exerting force on the particle [23]. Here, tachyon particles are avoided but it is not clear if and how the model can be verified.
@@
In this article, after observing that the more spontaneous relativistic generalization of the original de Broglie-Bohm theory is unsatisfactory as regards the continuity of the Quantum Potential with the non-relativistic case, we explore the strategy of introducing a new independent time parameter as the only hidden variable, whose relative motions are recognized to be instantaneous with respect to the traditional time. In this deterministic and explicitly non-local approach, a more general Zitterbewegung, as motion in the new time variable caused by the Quantum Potential, arises; then, it would be not directly observable, in accordance with recent beliefs [36, 37, 38]. The model is potentially capable to assign full physical reality to the Feynman paths and predicts, as non-standard fact, a modification of the uncertainty principle: by depending on the direction of motion, it would become anisotropic. This property could be observed in high energy accelerators.
@@ II Relativistic generalization issue for the quantum potential in case of constant momentum
Now we assume that the continuity equation (3), with $m\gamma$ instead of m, still holds. By employing $\vec{\nabla}\cdot\vec{v}=0$, we get just the expansion of $\frac{dR^{2}}{dt}$. Of course, we can refer it simply to R:
@@ III The two-time Bohmian model
In the first place it is therefore necessary to admit that, even in the case of a free particle, $\vec{p}$ is not constant. Indeed, inside the Heisenberg picture of the standard interpretation, Schrödinger and Dirac recognized that an omnidirectional velocity $\vec{v}_{i}$ of a sinusoidal motion of frequency $\frac{2m\gamma c{}^{2}}{\hbar}\,$ and amplitude about $\frac{\hbar}{2m\gamma c}\,$, known as Zitterbewegung [34, 35], must be superimposed on the actually observed velocity in average value. This motion has been much commented on, sometimes related to spin [41, 42, 43, 44] and Quantum Potential [45], but the debate on its exact interpretation, and even observability, does not seem yet resolved [36, 37, 38, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55]. Hopefully, to rediscover this peculiar movement inside a deterministic theory could clarify its nature. It can only be due to the Quantum Potential; but equation of motion (6) still seems inadequate: in fact it has been shown that putting $\frac{d\vec{p}_{i}}{dt}=-\vec{\nabla}V_{Q}$ leads to inconsistency in [1 + 1] dimensions [40].
Here we explore the effect of introducing a new independent time variable for the spatial coordinates of the particle. Two time physics has been considered in various contexts. An additional temporal parameter is considered e.g. in [56, 57, 58, 59, 60] and, with a different approach, in [61, 62, 63, 64] in order to more accurately explain different quantum properties or obtain supersymmetries. But in these theories the new time is not independent from the old one as we wish.
@@
$v_{cx}\equiv\frac{\partial X}{\partial t}\quad\quad v_{ix}\equiv\frac{\partial X}{\partial\tau}\,$ (16)
having so: $dX=v_{cx}dt+v_{ix}d\tau$. If we suppose that $v_{cx}$ and $v_{ix}$ depend, respectively, only on $t$ and $\tau$, we have:
@@
The intrinsic motion is not directly observable and, for the independence of $\tau$, is perceived as instantaneous with respect to time t. This explicitly non-local behavior has dramatic effects on the measurements of the observable quantities. As for the position, eq. (17) allows to the particle, observed in its trajectory as a function of t, to jump instantly from one point of space to another, in principle arbitrarily distant, through the time dimension $\tau$; and this must be generalized in every direction of space. Actually, Quantum Mechanics has accustomed us to even more surprising facts, not limited to the cases of an electron spreaded in an atomic orbital, quantum diffraction or entanglement, but really in any motion. Indeed, according to Feynman’s sum over histories interpretation (which is equivalent to the standard one) the physical event ”the particle (or photon) moves from A to B” is only correctly explained by admitting that it has traveled at once all the possible paths from A to B [65] [^1] Also interacting with every possible virtual particle of the vacuum, like experience excellently confirms..
For a particle in presence of a potential $V$ the wavefunction is then written:
@@
where the Quantum Potential does not appear; that is also true, naturally, for its non-relativistic approximation. This equation is therefore incompatible with the Bohm’s motion equation (6). On the other hand, if we take the total derivative over $t$ of eq. (4), we re-obtain the eq. (20), because $H-V_{Q}=E_{c}=const$. The model is thus in disagreement with the traditional non-relativistic de Broglie-Bohm theory. However, due to its setting, we will continue to call it Bohmian.
@@
As regards the omnidirectional intrinsic motions, it is spontaneous to consider that they are the direct cause of the quantum uncertainties. More, we assume that the Special Relativity does not hold for them. This condition is in agreement with the standard Zitterbewegung, where the mass of the particle does not undergo any relativistic increase - other than that one due to the classic velocity $\vec{v_{c}}$ despite it reaches the speed c. Being $m\gamma_{c}$ the mass involved in the generic intrinsic motion, we so have $\vec{p}_{i}=m\gamma_{c}\vec{v}_{i}$.
@@
$\frac{\partial R}{\partial\tau}-\vec{v_{i}}\cdot\vec{\nabla}R=0$ (22)
To justify it, we can observe that it is a form of continuity equation in which the addends subtract instead of adding up. This would remind the behavior of an antiparticle, although in the following we will not make any reference to antimatter and only use positive energy values. Since $\tau$-motion is not directly influenced by $V$, we can consider that $v_{i}$ only depends on $\tau$; consequently, eq. (22) holds for any function of any-order derivative of $R$ over $t$ or space. Hence we are admitting that it can be applied to $V_{Q}$, at least in case of conservative $V$.
Conservation of energy for a generic motion in $\tau$ reads:
$V_{Q}+\frac{1}{2}m\gamma_{c}v_{i}^{2}\equiv E_{q}$ (23)
@@
$m\gamma_{c}\frac{d\vec{v}_{i}}{d\tau}=-\vec{\nabla}V_{Q}$ (24)
The classic motion equation (20) and the generic $\tau$-motion (24) have so to replace the Bohm’s original one (6).
@@
    For $t$-motion and a generic $\tau$-motion the continuity eqs. (21) and (22), respectively, hold.
@@
Certainly, guidance equation $\vec{p}_{c}=\hbar\,Im\,(\frac{\vec{\nabla}\psi}{\psi})$ is obtained by taking the gradient of $\psi$ and applying $\vec{\nabla}S=\vec{p}_{c}$.
It should also be noted that, although each $\tau$-motion does not respect the Special Relativity, the maximum value of every $v_{i}$ is $c$: in fact it represents the quantum uncertainty on the homologous component of $v_{c}$, which cannot be greater than $c$.
The model therefore adds a Newtonian time to the usual four-dimensional $t$-spacetime. As we will see in the next sections, the motions in the new time variable are absolutely negligible for macroscopic objects. Conversely, when this is not the case, it happens that, even assuming to know the Quantum Potential with a good approximation, due to the absolute independence of the two times and the omnidirectionality of the $\tau$-motion, the function $\vec{r}(t,\tau)$ is unable to provide us the actual trajectory followed by the particle: it rather represents a three-dimensional object of $t$-spacetime. In particular, while $t$ remains constant, the particle can $\tau$-oscillate an arbitrary number of times in any arbitrary direction. However, this is an indeterminism linked to the non-existence of a specific function $\tau(t)$ and no longer to a conceptual dualism, as in the standard interpretation. In this model we still have a corpuscular and a wave point of view, the latter analogous to the standard one. However, there is no incompatibility between them, but the wave point of view is necessary due to the described constitutional insufficiency of the corpuscular description.
@@ IV Zitterbewegung in the new model
The classical motion equation (20) just informs us that $\vec{p}_{c}$ is constant. All equations found in section II - except (6) - are valid. Again, eq. (21) can be rewritten:
@@
So, $V_{Q}$ and H are waves that follow the particle in $t$-motion. Reasoning in the same way on eq. (22), we find instead that they are regressive waves in $\tau$-motion.
@@
$v_{ix}(\tau)\simeq\Delta X\,\frac{c}{\lambda}\,f\sqrt{2(f+2)})\,cos(\frac{c}{\lambda}\,f\sqrt{2(f+2)})\tau)$ (44)
where, according to point 1 of the theory, the amplitude $\Delta X$ represents the positional indeterminacy in $x$ of the particle. To get amplitude and angular frequency of the standard Zitterbewegung, i.e. $\frac{\hbar}{2m\gamma_{c}c}$ and $\frac{2m\gamma_{c}c{}^{2}}{\hbar}$, it is enough to set $\gamma_{c}\,f\sqrt{2(f+2)}=2$, and $v_{ix_{MAX}}=c$. However, by imposing only $v_{ix_{MAX}}=c$, we obtain $\Delta X=\frac{\lambda}{f\sqrt{2(f+2)}}$; as for f, we will simply assume that it is of the order of unity. Then, in the model under consideration, the uncertainty principle is:
@@
The amplitude of the $\tau$-oscillation - i.e. the quantum positional uncertainty - is inversely proportional to the rest mass. Then, for a particle of one picogram we find it about $15$ orders of magnitude smaller than for an electron. However, this is not the only reason that explains the evanescence of the $\tau$-motion for macroscopic bodies, as we will show in the next section.
@@
The total purely quantum energy $E_{q}\equiv V_{Q}+\frac{1}{2}m\gamma_{c}v_{ix}^{2}$ is independent of $x$:
@@
[…]
@@ IV.1 Zitterbewegung in arbitrary direction
In an arbitrary direction $\hat{s}$, which forms an angle $\vartheta$ with $\hat{v}_{c}$, eq. (25), for F=R, is written:
$\frac{\partial R}{\partial t}+{v_{cs}}\dot{R}=0$ (48)
where $v_{cs}=v_{c}\,cos\,\vartheta$ and the dot denots derivative over s. So eqs. (28) and (29) are generalized by:
$H=\sqrt{m^{2}\gamma_{c}^{2}c^{4}-\hbar{}^{2}\frac{c^{2}}{\gamma_{cs}^{2}}\frac{\ddot{R}}{R}}$ (49)
$v_{cs}\,\frac{\dot{R}}{R}\,(m\,\gamma_{c}\,c^{2}-H)+\frac{1}{2}\frac{\partial H}{\partial t}=0$ (50)
Placing: $\lambda^{2}\equiv\frac{\hbar{}^{2}}{m^{2}c^{2}\gamma_{c}^{2}\gamma_{cs}^{2}}$, we get back the same equations we have considered. Expressions of the Hamiltonian and Quantum Potential, given by (33) and (37), do not change, not even the amplitude of their oscillations. By imposing the motion equation:
$m\gamma_{c}\frac{d{v}_{is}}{d\tau}=-\dot{V}_{Q}$ (51)
we obtain an intrinsic harmonic motion described again by (43) and (44), but with the new value of $\lambda$. The consequence on the uncertainty principle - eqs. (45, 46) - is the presence of $\gamma_{cs}$ instead of $\gamma_{c}$ to divide. This gives rise to a spatial anisotropy of the uncertainty principle:
$\Delta s\cdot\Delta p_{s}\sim\frac{\hbar}{\gamma_{cs}}$ (52)
In particular, for $\hat{s}$ tending to be perpendicular to $\hat{v}_{c}$, $\gamma_{cs}\rightarrow 1$ and the standard uncertainty principle, independent of speed, is re-established. However, the case of exact perpendicularity is singular: for $v_{cs}=0$ eq. (50) implies $H=const\Rightarrow V_{Q}=const\Rightarrow\vec{v}_{is}=const$, instead of being oscillatory.
@@
For the non-relativistic ($v_{c}\ll c$) and $v_{c}\rightarrow 0$ cases, it is sufficient to approximate $\gamma_{c}$ in the previous equations (respectively, with $1+\frac{v_{c}^{2}}{2c^{2}}$ and 1), obtaining still high frequency oscillatory motions in $\tau$. These results cannot be obtained starting from the standard Schrödinger equation because from it the non-relativistic approximation of eq. (50) cannot be deduced. This happens since the explicit dependence of $R$ on $t$ causes the Klein-Gordon equation to no longer tend to the Schrödinger equation.
In $v_{c}=0$ there is another singularity; in fact, by eq. (50) we get again $\vec{v}_{is}=const$, contrary to the case $v_{c}\rightarrow 0$.

### v16

@@ I Introduction
    Rejecting the decomposition of the wavefunction into norm and phase, and interpreting it as a real quantum field exerting force on the particle [23]. Here, tachyon particles are avoided but it is not clear if and how the model can be verified.
@@
In this article, after observing that the more spontaneous relativistic generalization of the original de Broglie-Bohm theory is unsatisfactory as regards the continuity of the Quantum Potential with the non-relativistic case, we explore the strategy of introducing a new independent time parameter as unique hidden variable, whose relative motions are recognized to be instantaneous with respect to the traditional time. In this deterministic and explicitly non-local approach, a more general Zitterbewegung, as motion in the new time variable caused by the Quantum Potential, arises; then, it would be not directly observable, in accordance with recent beliefs [36, 37, 38]. The model is potentially capable to assign full physical reality to the Feynman paths and predicts, as non-standard fact, a modification of the uncertainty principle: by depending on the direction of motion, it would become anisotropic. This property could be observed in high energy accelerators.
@@ II Relativistic generalization issue for the quantum potential in case of constant momentum
Now we assume that eq. (3), with $m\gamma$ instead of m, still holds (we will justify later this continuity equation). By employing $\vec{\nabla}\cdot\vec{v}=0$, we get just the expansion of $\frac{dR^{2}}{dt}$. Of course, we can refer it simply to R:
@@ III The two-time Bohmian model
In the first place it is therefore necessary to admit that, even in the case of a free particle, $\vec{p}$ is not constant. Indeed, inside the Heisenberg picture of the standard interpretation, Schrödinger and Dirac recognized that an omnidirectional velocity - which we will call $v_{i}$, with $i$ standing for intrinsic - of a sinusoidal motion of frequency $\frac{2m\gamma c{}^{2}}{\hbar}\,$ and amplitude about $\frac{\hbar}{2m\gamma c}\,$, known as Zitterbewegung, must be superimposed on the actually observed velocity in average value [34, 35]. This motion has been much commented on, sometimes related to spin [41, 42, 43, 44] and Quantum Potential [45], but the debate on its exact interpretation, and even observability, does not seem yet resolved [36, 37, 38, 46, 47, 48, 49, 50, 51, 52, 53, 54, 55]. Hopefully, to rediscover this peculiar movement inside a deterministic theory could clarify its nature. It can only be due to the Quantum Potential; but equation of motion (6) still seems inadequate: in fact it has been shown that putting $\frac{dp_{i}}{dt}=-\mid\vec{\nabla}V_{Q}\mid$ leads to inconsistency in [1 + 1] dimensions [40].
Here we explore the effect of introducing a new independent time variable for the spatial coordinates of the particle. This outcome certainly deserves some reasons. The Feynman’s sum over histories interpretation of Quantum Mechanics (compatible with the standard one) actually operates as if matter could reside simultaneously in multiple places in space. If we assume this ”as if” really be a ”so it is”, then the problem arises of explaining how this can happen. An object can occupy different positions in space thanks to movement, which is parameterized by the usual time. If we imagine that there exists an extra time independent of the latter - that is, free to flow while the usual time is frozen -, then there may be a hidden movement capable of assigning to the object different positions, which so will appear simultaneous with respect to the usual time. It is of fundamental importance to keep in mind that this simultaneity is totally different from the one that can exist between two $t$-events, which is considered (and rightly relativised) in Special Relativity. Here the recognized hidden motion occurs with respect to a new and independent temporal parameter and therefore, as long as the object does not interact with something to constitute a $t$-event, it must be considered as simultaneous in any reference system. Section V will be dedicated to the in-depth analysis of this phenomenology.
Two time physics has been considered in various contexts. An additional temporal parameter is considered e.g. in [56, 57, 58, 59, 60] and, with a different approach, in [61, 62, 63, 64] in order to more accurately explain different quantum properties or obtain supersymmetries. But in these theories the new time is not independent from the old one as we wish.
@@
$v_{c_{x}}\equiv\frac{\partial X}{\partial t}\quad\quad v_{i_{x}}\equiv\frac{\partial X}{\partial\tau}\,$ (16)
having so: $dX=v_{c_{x}}dt+v_{i_{x}}d\tau$. If we suppose that $v_{c_{x}}$ and $v_{i_{x}}$ depend, respectively, only on $t$ and $\tau$, we have:
@@
The intrinsic motion is not directly observable and, for the independence of $\tau$, is perceived as instantaneous with respect to time t. This explicitly non-local behavior has dramatic effects on the measurements of the observable quantities. As for the position, eq. (17) allows to the particle, observed in its trajectory as a function of t, to jump instantly from one point of space to another, in principle arbitrarily distant, through the independent time dimension $\tau$; and this has to be generalized in every direction of space. Actually, Quantum Mechanics has accustomed us to even more surprising facts, not limited to the cases of an electron spreaded in an atomic orbital, quantum diffraction or entanglement, but really in any motion. Indeed, according to Feynman’s sum over histories interpretation (which is equivalent to the standard one) the physical event ”the particle (or photon) moves from A to B” is only correctly explained by admitting that it has traveled at once all the possible paths from A to B, also interacting with every possible virtual particle of the vacuum [65]. And this is excellently confirmed by experience.
Let’s continue now to specify the dynamics of the model keeping in mind that, due to its omnidirectionality, the intrinsic velocity cannot be treated as a vector.
For the generic particle wavefunction:
@@
we admit, justifying it later, a two-time action $S=S_{t}+S_{\tau}$ so defined:
[…]
@@
where the Quantum Potential does not appear; that is also true, naturally, for its non-relativistic approximation. This equation is therefore incompatible with the Bohm’s motion equation (6). On the other hand, if we take the total derivative over $t$ of eq. (4), we re-obtain the eq. (20), because $H-V_{Q}=E_{c}=const$. The model is thus in disagreement with the traditional non-relativistic de Broglie-Bohm theory. Nevertheless, due to its setting, we will continue to call it Bohmian.
@@
It expresses the conservation of the probability of finding the particle in a unit of volume and its validity, undisputed in the non-relativistic case, must extend to the relativistic one. In fact, this probability is independent of the relativistic mass of the particle and could depend only on its Lorentzian dimensional contraction, which is neglected when treating the particle as a material point. On the other hand, the identical equation is valid in Relativity for the electric charge density.
As regards the intrinsic motions, it is spontaneous to consider that they are the direct cause of the quantum uncertainties. However, if this is assumed, then we must exclude that Special Relativity holds for them. We will conclude this in the next section, studying the case of the free particle. For now, we just observe that this condition is in agreement with the standard Zitterbewegung, where the mass of the particle does not undergo any relativistic increase - other than that one due to the classic velocity $\vec{v_{c}}$ despite it reaches the speed c.
If this is assumed, we have: $p_{i_{s}}=m\gamma_{c}v_{i_{s}}$, which justifies the previous definition of $S_{\tau}$.
@@
$\frac{\partial R}{\partial\tau}-v_{i_{s}}{\frac{dR}{ds}}=0$ (22)
for any $\hat{s}$. It establishes that $R$ is a regressive wave, which moves in a specular manner with respect to the particle. We will provide a justification of this key equation in the Appendix.
Since $\tau$-motions are not directly influenced by $V$, we can consider that $v_{i_{s}}$ only depends on $\tau$; consequently, eq. (22) holds for any function of any-order derivative of $R$ over $t$ or space. Hence we are admitting that it can be applied to $V_{Q}$, at least in case of conservative $V$.
Conservation of energy for the generic $\tau$-motion reads:
$V_{Q}+\frac{1}{2}m\gamma_{c}v_{i_{s}}^{2}\equiv E_{q_{s}}$ (23)
@@
$m\gamma_{c}\frac{dv_{i_{s}}}{d\tau}=-\frac{dV_{Q}}{ds}$ (24)
for every $\hat{s}$. Eqs. (20) and (24) have so to replace the Bohm’ motion eq. (6).
@@
    For $t$-motion and each $\tau$-motion the continuity eq. (21) and the wave eq. (22), respectively, hold.
@@
Certainly, guidance equation $\vec{p}_{c}=\hbar\,Im\,(\frac{\vec{\nabla}\psi}{\psi})$ is still obtained by taking the gradient of $\psi$ and applying $\vec{\nabla}S=\vec{p}_{c}$.
It should also be noted that, although each $\tau$-motion does not respect the Special Relativity, the maximum value of $v_{i_{s}}$ is $c$: in fact it represents the quantum uncertainty on the generic component of $\vec{v_{c}}$, which cannot be greater than $c$.
The model therefore adds a Newtonian time to the usual four-dimensional $t$-spacetime. As we will see in the next sections, the motions in the new time variable are absolutely negligible for macroscopic objects. Conversely, when this is not the case, it happens that, even assuming to know the Quantum Potential with a good approximation, due to the absolute independence of the two times and the omnidirectionality of the $\tau$-motions, the function $\vec{r}(t,\tau)$ is unable to provide us the actual trajectory followed by the particle: it rather represents a three-dimensional object of $t$-spacetime. In particular, while $t$ remains constant, we will find that the particle can $\tau$-oscillate an arbitrary number of times in any arbitrary direction. However, this is an indeterminism linked to the non-existence of a specific function $\tau(t)$ and no longer to a conceptual dualism, as in the standard interpretation. In this model we still have a corpuscular and a wave point of view, the latter analogous to the standard one. However, there is no incompatibility between them, but the wave point of view is necessary due to the described constitutional insufficiency of the corpuscular description.
In this section we have outlined the two-time model, highlighting its basic properties. In the next one we will study the free particle.
@@ IV Zitterbewegung in the new model
The classical motion equation (20) just informs us that $\vec{p}_{c}$ is constant. All equations (7) through (14) found in section II are valid. Again, eq. (21) can be rewritten:
@@
So, $V_{Q}$ and H are progressive waves, i.e. they follow the particle in its $t$-motion.
@@
$v_{i_{x}}(\tau)\simeq\Delta X\,\frac{c}{\lambda}\,f\sqrt{2(f+2)})\,cos(\frac{c}{\lambda}\,f\sqrt{2(f+2)})\tau)$ (44)
where, according to point 1 of the theory, the amplitude $\Delta X$ represents the positional indeterminacy in $x$ of the particle. To get amplitude and angular frequency of the standard Zitterbewegung, i.e. $\frac{\hbar}{2m\gamma_{c}c}$ and $\frac{2m\gamma_{c}c{}^{2}}{\hbar}$, it is enough to set $\gamma_{c}\,f\sqrt{2(f+2)}=2$, and $v_{i_{x_{MAX}}}=c$. However, by imposing only $v_{i_{x_{MAX}}}=c$, we obtain $\Delta X=\frac{\lambda}{f\sqrt{2(f+2)}}$; as for f, we will simply assume that it is of the order of unity. Then, in the model under consideration, the uncertainty principle is:
@@
Before continuing the analysis it is opportune to comment on what we would have found if we had assumed the Special Relativity be valid for $\tau$-motion. The study of a relativistic oscillator [66], provides the following result in our case:
$\Delta X=\frac{c\sqrt{2}\sqrt{1-\sqrt{1-\beta^{2}}}}{\omega_{0}(1-\beta^{2})^{\frac{1}{4}}}$ (46)
where $\omega_{0}$ is the non-relativistic pulsation and $\beta=\frac{v_{i_{x_{MAX}}}}{c}$. But this is incompatible with our model: by imposing $v_{i_{x_{MAX}}}=c$, we obtain $DX\rightarrow\infty$, therefore not being able to represent the minimum positional uncertainty.
The amplitude of the $\tau$-oscillation, i.e. the quantum positional uncertainty, is inversely proportional to the rest mass. Then, for a particle of one picogram we find it about $15$ orders of magnitude smaller than for an electron. However, this is not the only reason that explains the evanescence of the $\tau$-motions for macroscopic bodies, as we will show in the next section.
@@
The purely quantum total energy $E_{q_{x}}\equiv V_{Q}+\frac{1}{2}m\gamma_{c}v_{i_{x}}^{2}$ is independent of $x$:
@@
It is noteworthy that Zitterbewegung in $\tau$-time is able to justify the physical origin of the phase to be associated with each path in Feynman’s formulation of Quantum Mechanics. The approach has been specified in [67], through a particle’s oscillation forward and backward in t-time. In this model, that type of motion arises spontaneously: it is the found Zitterbewegung in $\tau$, represented by the wavefunction extra phase factor $e^{\frac{i}{\hbar}S_{\tau}}$.
We have so found the Zitterbewegung in direction of motion; in the following three subsections we generalize and complete its study.
@@ IV.1 Zitterbewegung in arbitrary direction
With respect to an arbitrary spatial direction $\hat{s}\neq\hat{v}_{c}$, let’s consider the polar coordinate system ($s$, $\theta$, $\phi$), where $\theta\neq 0$ is the angle between $\hat{s}$ and $\hat{v}_{c}$ and $\phi$ the azimuth (see next figure).
Having as our goal to find the $\tau$-motion along $s$ by eq. (24), $\theta$ and $\phi$ must be considered constant, so the polar expressions of $\vec{\nabla}R$ and $\nabla^{2}R$ reduce to:
$\vec{\nabla}R=\frac{dR}{ds}\hat{s}\quad\quad\nabla^{2}R=\frac{2}{s}\frac{dR}{ds}+\frac{d^{2}R}{ds^{2}}$ (49)
substituting the first one into eq. (25) and differentiating on $t$ gives:
$\frac{\partial^{2}R}{\partial t^{2}}-{v_{c_{s}}^{2}}R^{\prime\prime}=0$ (50)
where $v_{c_{s}}=v_{c}\ cos\theta$ and the apostrophe indicates derivative over $s$. Substituting eq. (50) and the second of eqs. (49) into eq. (11) and rooting is found:
$H=\sqrt{m^{2}\gamma_{c}^{2}c^{4}-\hbar^{2}\frac{c^{2}}{\gamma_{c_{s}}^{2}}\frac{R^{\prime\prime}}{R}+\frac{2\hbar^{2}}{sR}R^{\prime}}\ ,\ \gamma_{c_{s}}\neq\gamma_{c}$ (51)
while eq. (14) is generalized by:
$v_{c_{s}}\,\frac{R^{\prime}}{R}\,(m\,\gamma_{c}\,c^{2}-H)+\frac{1}{2}\frac{\partial H}{\partial t}=0$ (52)
Now, remembering that, in order to integrate eq. (38), we had to limit ourselves to small neighborhoods of the maximum $R_{M}$, we can neglect the last addend of eq. (51) and thus lead back to the already studied eqs. (28) and (29). Placing: $\lambda^{2}\equiv\frac{\hbar{}^{2}}{m^{2}c^{2}\gamma_{c}^{2}\gamma_{c_{s}}^{2}}$, we so get back the same equations found before. Expressions of the Hamiltonian and Quantum Potential, given by (33) and (37), do not change, not even the amplitude of their oscillations. By imposing the $\tau$-motion equation, we obtain an intrinsic harmonic motion described again by (43) and (44), but with the new value of $\lambda$.
The different value of $\lambda$ therefore causes a dependence of the oscillation frequency on the spatial direction: this is precisely the relativistic anisotropy $v_{i_{s}}\neq v_{i_{u}}$ that we anticipated in section III. Remarkably, this is reflected in a spatial anisotropy of the uncertainty principle, where we have the presence of $\gamma_{c_{s}}$ instead of $\gamma_{c}$ to divide:
$\Delta s\cdot\Delta p_{s}\sim\frac{\hbar}{\gamma_{c_{s}}}$ (53)
In particular, for $\hat{s}$ tending to be perpendicular to $\hat{v}_{c}$, $\gamma_{c_{s}}\rightarrow 1$ and the standard uncertainty principle, independent of speed, is re-established. However, the case of exact perpendicularity is singular: for $v_{c_{s}}=0$ eq. (52) implies $H=const\Rightarrow V_{Q}=const\Rightarrow{v}_{i_{s}}=const$, instead of being oscillatory.
@@
For the non-relativistic ($v_{c}\ll c$) and $v_{c}\rightarrow 0$ cases, it is sufficient to approximate $\gamma_{c}$ in the previous equations (respectively, with $1+\frac{v_{c}^{2}}{2c^{2}}$ and 1), obtaining still high frequency oscillatory motions in $\tau$. These results cannot be obtained starting from the standard Schrödinger equation because from it the non-relativistic approximation of eq. (52) cannot be deduced. This happens since the explicit dependence of $R$ on $t$, causes the Klein-Gordon equation to no longer tend to the Schrödinger equation.
In $v_{c}=0$ there is another singularity: as before, we get by eq. (52) ${v}_{i_{s}}=const$ for any $\hat{s}$, contrary to the case $v_{c}\rightarrow 0$.